Assume the economy as reflected in the equilibrium point of AD/AS is at full employment GDP at an output of $17 trillion

  1. Consumer confidence has increased with news of large stock market gains. As a result, $ 1 trillion dollars of spending increases from consumers. What happens initially to AD and what is the level of output now?
  2. Assuming a mpc of 0.8, what will happen to real GDP after all rounds of spending have been completed? Show your calculation to receive full points.
  3. In two sentences, explain there is a multiplier effect on the economy
Answer #1

a) With increase in consumer confidence from news of large stock market gains, AD curve shifts to the right. As a result, level of output must increase.

b)MPC=0.8

Government spending multiplier = 1/(1-MPC) = 1/(1-0.8) = 1/0.2 =5

As a result, with $1 trillion increase in government spending, output Y increases by $1 trillion*5 = $5 trillion .

With any new injection (government spending) in the economy, output increases by a greater amount. This is the multiplier effect.